The acceleration a of a particle varies with time t as a=bt+c where b and c are constants.What will be the velocity of the particle which starts from rest after the time t?
 

Grade:11

1 Answers

sreehari
6 years ago
\frac{\mathrm{d} v}{\mathrm{d} t}= a
where v is velocity and a is acceleration.
Hence,
v(t)-v(0) = \int_{0}^{t} adt
where v(t) is velocity at time t and v(0) is velocity at time 0.
In this question
 
v(t)=\frac{bt^{2}}{2} + ct